Trace the curve x^3 − 6x^2 + 11x − y = 6.
1
Expert's answer
2020-12-21T19:33:46-0500

Equation of the curve is "y = x^3-6x^2+11x-6"


This equation can be written as,

"y = (x-1)(x-2)(x-3)"

When x=0, y=-6.


So, I can directly find the x where y=0,

It will give me, "x= 1,2,3"

So I will have intervals as "(-\infty, 1], [1,2],[2,3],[3,\infty)"


In interval "(-\infty, 1]", If I put any value from "-\infty" to 1, value of the y will increase.

In interval "[3, \infty)" , on increasing value of x, y will increase.


Now, "\frac{dy}{dx} = 3x^2-12x+11"

Putting "\frac{dy}{dx} =0," We get, "x = 1.423,2.577"


"\frac{d^2 y}{dx^2} = 6x-12"


At x=1.423, "\frac{d^2 y}{dx^2} = 6(1.423)-12 < 0", Hence it is point of maxima.


At x=2.577, "\frac{d^2 y}{dx^2} = 6x-12 > 0" , So it is point of minimum.


Using all these properties, We find that curve will look like,